OpenStudy (anonymous):

Need help with system: -4x-2y=14 -10x+7y=-25

OpenStudy (saifoo.khan):

have u tried it? where are u stuck?

OpenStudy (anonymous):

i need to know what to multiply what to begin elimination

OpenStudy (saifoo.khan):

first decide which one do u want to eliminate?

OpenStudy (anonymous):

-10x+7y=-25

OpenStudy (saifoo.khan):

No, which variable.. x or y/

OpenStudy (anonymous):

y

OpenStudy (saifoo.khan):

Alright, now if u make both 14, you can easily cancel them out. to that you have to multiply both of the equations... equation 1 with (7) -4x-2y=14 equation 2 with 2. -10x+7y=-25

OpenStudy (anonymous):

ooooooooooooooooooooooooooooooooooooooooooooooooooh...yeah i knew that...

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

this will give you, -28x -14y = 98 -20x + 14y = -50

OpenStudy (saifoo.khan):

-48x = 48 x = -1

OpenStudy (saifoo.khan):

And so on..... :)

OpenStudy (anonymous):

yeah i knew that.............:(....................thanks